Find the sum of the first 10 terms of the geometric series 4-8+16-32+64
Answer by fractalier(6550)

There is a closed form solution for this, but without that at my finger tips I would suggest looking at pairs...
4-8 = -4
16-32 = -16
64-128 = -64
256-512 = -256
1024-2048 = -1024
Add 'em up and get
-1364